Create a real-life situation that fits into the equation (x + 4)(x - 7) = 0 and express the situation as the same equation. I don't know how to work this problem backwards, help me please.
Answer by Mathtut(3670)

Create a real-life situation that fits into the equation (x + 4)(x - 7) = 0 and express the situation as the same equation:
:
This would be one example:
:
The area of a rectangle is 28 sq ft,
where the width is 3 ft shorter than the Length.
:
width: W = (L - 3)
:
Area = L * (L-3) = 28
:
L^2 - 3L - 28 = 0
(L + 4)(L - 7) = 0
L = 7